A gas stattion sells 4820 gallons of gas on a day when they charge $1.35 per gallon, and they sell 3880 gallons on a day when they charge $1.40. Find the equation for the line whose graph depicts this data.
Answer by alicealc(293)

price = x
number of gallons sold = y
y = ax + b
(1) 4820 = 1.35a + b
(2) 3880 = 1.40a + b
--------------------- (-)
940 = -0.05a
a = 940/-0.05
a = 94000/-5 = -18800

4820 = 1.35a + b
4820 = 1.35*(-18800) + b
4820 = -25380 + b
b = 4820 + 25380 = 30200

so, the line equation for the problem is: y = -18800x + 30200